OpenStudy (anonymous):

Would the equation 3x^2+x-5=0 have two complex roots?

OpenStudy (anonymous):

Factor out 3x^2-x-5 (3x+5)(x-2) Set each equal to 0 3x+5=0 x-2=0 solve X=-5/3 and x=2

OpenStudy (anonymous):

no the value of \[b^2-4ac>0\] . hence no complex roots...it will have two distinct real roots

OpenStudy (anonymous):

Could you show me how you got that. I guess I just don't understand it very well. Please :)

OpenStudy (anonymous):

it is actually kinda formula. \[\Delta=b^2-4ac \] it is called the discriminant if \[\Delta>0 ( we \space have \space two \space real \space distinct \space roots)\] \[\Delta=0 ( we \space have \space two \space real \space but \space equal \space roots)\] \[\Delta<0 ( we \space have \space two \space complex \space roots)\]

OpenStudy (mrnood):

@answers4you Factor out 3x^2-x-5 (3x+5)(x-2) = 3x^2-x -10 therefore not a correct factorisation